# Breaking the Trend: How to Avoid Cherry-Picked Signals

### v10

@@ Abstract
Our empirical results show an impressive fit with the pretty complex theoretical Sharpe formula of a Trend following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014). That empirical fit convinces us that a mean-reversion process with only one time scale is enough to model, in a pretty precise way, the reality of the trend-following mechanism at the average scale of CTAs and as a consequence, using only one simple $\mathrm{EMA}$, appears optimal to capture the trend. As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
December 24, 2025
@@ 2 Literature
The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as a signal for each underlying instrument (most often individual indicator gives either a long or a short position [^1] Sepp and Lucic (2025) classified the methodologies for sizing positions based on signals into three categories: (1) European trend following, which assumes continuous position sizes proportional to signal strength — an approach used by many large European CTA managers; (2) Time Series Momentum (Moskowitz et al (2012); Hurst et al (2013); Baltas (2015)), a method more commonly used in academic studies, where positions depend on the sign of the momentum; and (3) American trend following, which originates from the early adopters of CTA strategies.). Then, a risk management process is applied to the signals to ensure that the portfolio is diversified enough and targets a constant risk. One simple solution is to impose an equal conditional risk in each asset class, as the SG Trend Indicator does and to size the portfolio to target a constant volatility. Another solution is to use the correlation matrix and apply mathematical optimization to maximize the reward/risk ratio. The well-known Markowitz solution set positions as linearly depending on the signals. The linear dependence is derived simply througth the normalization of the signals by the inverse of the correlation matrix between returns, but the empirical results are not appealing. The Agnostic Risk Parity ($\mathrm{ARP}$), introduced by Benichou at al (2017), normalizes the signals through the inverse of the square root of the correlation matrix. Benichou at al (2017) proposed this portfolio because it was rotationally invariant. Since this concept is primarily understood by physicists, their portfolio is not yet widely accepted in the finance community, with only a few citations, despite posing a serious challenge to Markowitz. For example, Benveniste at al (2024) explained, for the celebration of the 50th year of the Journal of Portfolio Management, the myth surrounding the alleged inadaptation of mean–variance optimization to the real world, and how efficient mean-variance optimization works without any conflict with empirical measurements, as long as the correlation matrix is properly modeled, yet neither referenced Benichou at al (2017) nor addressed the impact of uncertainty in mean within Markowitz formula. The literature has long acknowledged the sensitivity of the Markowitz formula to parameter uncertainty, without recognizing that this uncertainty can actually alter the form of the formula itself [^2] A dedicated review of the literature is provided in the appendix..
@@
In reality, one hidden assumption of Markowitz optimization is that the expected means of returns from the signals are known and certain, but this assumption is incorrect. Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal in the context of trend following, where expected means are not certain, are estimated with some errors and do not break rotational invariance as biases introduced by risk premia would. This optimality holds when the correlation matrix between the trends, signals or uncertainty of mean is different from the correlation matrix between returns but is a kind of random matrix where a dominant factor of this correlation matrix is very dominant and diffuses randomly. Valeyre (2024) suggested an interacting agents model to justify such an hypothesis.
@@
There are many different technical indicators used as proxies for expected means in trend-following, but all of them are based on past returns, which are supposed to help predict future ones, while inevitably containing noise and uncertainty. Most of the popular indicators used are based on moving averages of past prices. The most popular is the Simple Moving Average ($\mathrm{SMA}$), while less commonly used types include the Linear Moving Average and the Exponential Moving Average on returns ($\mathrm{EMA}$). Each moving average is computed using an averaging window of a particular size. Trend-following indicators can also be based on a combination of moving averages, such as crossovers (one with a short window size and another with a long window size). The switch is determined when the short-term moving average crosses the long-term moving average. As an example, The Societe Generale Trend Indicator, which is a reference among CTAs, is detemined only by the crossover based on a simple average based 20 and 120 business day parameters. Momentum ($MOM$) is also a very popular indicator and is defined as the difference between two prices. Hurst et al (2017) analysed the performance simulated on the last century using a mixture of 1 month, 3 month and 1 year momentum. Lamperiere et al (2014) analysed the performance simulated on the last 2 centuries using a 50 business day EMA signal [^3] They were ambiguous and only wrote “exponential moving average […] with a decay rate equal to $n$ months.” The parameter $n$ is then interpreted as the half-life, defined as the lookback period over which the cumulative weight of past observations reaches $50\%$. This corresponds approximately to $\frac{\ln(2)}{\eta}$, where $\eta$ is defined later in Eq. (6).. Gmür et al (2025) specifies that a half-life of 60-70 business days for an $\mathrm{EMA}$, corresponding to approximately one calendar quarter, best replicates the return and risk characteristics of trend-following CTA industry benchmarks such as the SG Trend Index, which invests roughly on the 10 largest systematic CTA funds with trend-following behavior, making it a robust bottom-up representation of that style. The Bollinger Band ($\mathrm{BB}$) is very popular, but non continuous with only 3 outputs 0 and 1 or -1 when price is oustide a band. The Moving Average Convergence/Divergence ($\mathrm{MACD}$) uses a combination of three EMAs to capture mean reversion at both short and long-term scales while identifying trends in the long term.
@@
We can describe some of these indicators through their sensitivity to past daily returns as introduced by Zakamuli and Giner (2020). Fig.1 displays some cases. It can be observed that the Exponential Moving Average ($\mathrm{EMA}$) exhibits a sensitivity that declines exponentially, while the crossover resembles a hat shape. Sensitivity is low for very recent past returns, high for a moderately distant past, and negligible for a long past. The shape of the crossover can be adapted if price behavior exhibits mean-reverting tendencies on very short time scales, as seen in the case of individual stocks with low liquidity. From an academic point of view, almost all these trend-following rules are ad hoc and lack theoretical justification for their optimality. Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33 business day EMA 444i.e $\eta=\frac{1}{33}$ in Eq.6 should be the optimal EMA to replicate the mixture of 1 month, 3 month and 1 year momentum used by Hurst et al (2017) while a 43 business day EMA should be the optimal one for the SG trend Indicator crossover). Zakamuli and Giner (2020) also reminds us that it is well known that if returns follow an autoregressive process, the best predictor has the same functional form as the autoregressive process being predicted. Therefore, the expected structure of the autoregressive process should provide a good proxy for the best predictor.
@@
Grebenkov and Serror (2014) proved that, in the more complex yet more insightful case where portfolio positions depend linearly on the signal [^5] Grebenkov and Serror (2014) most likely introduced into the literature the ”linear trading rule” whose position depends linearly on the indicator, rather than a purely binary trading rule, which had been the standard approach most likely because the standard approach focuses on a one-dimensional problem and on a strategy that targets a constant volatility. We were unable to find any prior published articles except Brandt and Santa Clara (2006) described in the appendix, despite consulting ChatGPT, Google Scholar, and several professors in finance. However, as we are not specialist in this specific area of the literature, it is possible that we may have missed some references. Grebenkov and Serror (2014) introduced that linear methods in the literature because I was a colleague of both and I used that method to ensure that my statistical arbitrage trading program managing more than 500 dimensions or more than 500 single stocks at John Locke Investments maintained the property of “rotational invariance.” A linear trading rule was indeed necessary to have that rotational invariance property and design consistent strategies capable of performing trend following either on an individual stock, or a basket of stocks, or on the eigenvectors of the correlation matrix of individual stocks, rather than independently applying trend following to each stock in isolation. , the $\mathrm{EMA}$ with a larger decay parameter than the decay parameter of the autoregressive coefficients of the return process yields the optimal portfolio. This finding is in stark disagreement with Zakamuli and Giner (2024), simply because Grebenkov and Serror (2014) solved a different, yet superior, optimization problem. Indeed, assuming that positions depend linearly on the signal appears much more natural and is not as suboptimal as merely assuming a binary choice between a long or short position with the same level of risk withtout the possibility of implementing “money management”. Most important the version of Grebenkov and Serror (2014) is in line with the ”rotational invariance” property, the Markowitz optimization and Agnostic risk parity by Benichou at al (2017) where positions are linearly depending on the signals. The linear trading rule can be extended from the one-dimensional case to the multi-dimensional setting, allowing the incorporation of the correlation matrix to better adjust portfolio positions in order to target the maximum Sharpe ratio at the portfolio level.
@@
Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant666the formula proposed by Acar (2003), which depends on the correlation between the indicator and returns, leads-when expressed using the Grebenkov process-to an extremely cumbersome formula linking the parameters of the return process and the relaxation time of the EMA, which can hardly be described as objectively elegant (description of the different formulas and the derivations when assuming the process of Grebenkov are available in the Appendix B). Furthermore, Zakamulin and Giner (2022) and Zakamuli and Giner (2024) are both based on the formula from Acar (2003) and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$. That formula is key in our main objective as we want to validate it using empirical measurements.
@@
Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community 777the formula in Acar (2003); Zakamulin and Giner (2022); Zakamuli and Giner (2024) holds only with binary trading rules (proof in Appendix B.3, originaly derived by Firoozye and Koshiyama (2020)). Following Grebenkov and Serror (2014), Ferreira et al. (2018) published a similar formula for a linear trading rule, but in the context of a more general process with a simple moving average and not an EMA. Ferreira et al. (2018) did not mention Acar (2003) and derived his results through his own way. Later, after Grebenkov and Serror (2014), Firoozye and Koshiyama (2020) extended the formula of Acar (2003) to linear trading rules, yet these developments were ignored by Zakamulin and Giner (2022) and Zakamuli and Giner (2024), who continued to focuse exclusively on binary rules while ignoring Grebenkov and Serror (2014), the literature continued to focus on one-dimensional case and the wrong optimization problem—one where positions do not depend linearly on the signals. Dai at al (2016) addressed this issue by imposing a constraint that allowed only two possible positions (long or zero). Similarly, Nguyen at al (2014b) determined the optimal threshold for triggering either a long or short position.
@@
Our first and main objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the “Agnostic Risk Portfolio” ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$. This methodology ultimately extends the idea of Ferreira et al. (2018) by focusing on an AR(1) model to capture trends and by aggregating all markets through projection onto the $ARP$, rather than analyzing each market individually — which carries the risk of excessively noisy measurements and requires fitting complex and distinct ARIMA models.
@@ 3 Theory from Grebenkov and Serror (2014)
This classical one-dimensional model can be easily extended to multiple dimensions. In the multivariate case, two covariance matrices are required to account for correlations. The first matrix captures the correlations between the random components of the different underlyings, while the second matrix reflects the correlations among the AR(1) processes associated with each underlying. The model is formally described in Eq. 1, using the same notation as in Grebenkov and Serror (2014). In that equation the returns $r_{i,t}$ of the $i$th instrument among the N is described in two components: The first $\epsilon_{i,t}$ is a noise which is not autocorrelated while the second $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ is an AR(1) process which is autocorrelated where $\xi_{i,k}$ is a noise which is not autocorrelated. Ultimately Eq.1 describes the returns of a diffusive process with a positive autocorrelation model for returns $r_{i,t}$: The weights $\left(1-\lambda\right)^{t-1-k}$ are declining exponentialy. The parameter $\lambda$ describes the inverse of the relaxation time of the mean reversion AR(1) process describing the short bias, and $\beta$ represents the strength of that short stochastic bias in the returns. The noises $\xi_{i,k}$ are independent in the $k$ axis but may be correlated with $i$ axis (Valeyre (2024) introduces specific forms of that matrix, leading to particular optimal portfolios). The noises $\epsilon_{i,t}$ are independent in the $t$ axis but may be correlated with $i$ axis with a very close correlation matrix between returns. Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ and it follows an auto-regressive model of order 1, AR(1).
@@
Grebenkov and Serror (2014) solved the correct theoretical optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals when assuming that returns are well described by Eq.1. They considered only the one-dimensional case, but later, Grebenkov and Serror (2015) and Valeyre (2024) extended the analysis to the multi-dimensional setting.
@@
Instead of focusing on the best predictor or on the optimal portfolio with suboptimal constraints—such as the possibility of being either long with the same risk or flat, which is unfortunately common in the literature—Grebenkov and Serror (2014) assumes that positions should better depend linearly on the signals [^9] They considered only the EMA as the signal, which was a good choice, as Zakamuli and Giner (2024) showed that the best predictor has the same functional form as the autoregressive process being predicted. That is, in the case of Eq. 1, the optimal predictor should be an EMA, which they indeed used, which is a natural approach which is in agreement with the Markowitz solution. They then determine the optimal signal that maximizes the Sharpe ratio. Thus, Grebenkov and Serror (2014) introduced substantial novelty, even though it is not yet widely considered by most academics specializing in trend-following.
@@
In detail, they derived an explicit formula, in Eq. 3, the theoretical Sharpe ratio, ${\mathbf{S}}\left(\eta\right)$, when assuming returns are generated through their autocorrelation model in Eq. 1, as a function of $\eta$ (the smoothing parameter of the trend indicator $EMA$ which is defined mathematically later in Eq.6 ) , $\lambda$ (the inverse of the time scale of the mean reversion process that describes the trends in Eq. 1) and $\beta$ (the weight of the trend component in the returns Eq. 1) or its normalized version $\beta_{0}$ while also considering the inclusion of trading costs, denoted as $\theta$. Equation Eq.3 is in fact a remarkably elegant formula. As expected, the theoretical Sharpe ratio depends on $\beta$ (or $\beta_{0}$ its normalized version), the weight of the trend in the returns, in an approximately proportional manner when $\beta_{0}^{2}$ is significantly larger than $\frac{1}{\lambda}$ and when $\eta$ remains close to $\lambda$. In that case, the Sharpe ratio from Eq.3 can be approximated by $\beta_{0}\sqrt{\frac{1}{2}}$ or $\beta_{0}\sqrt{\frac{255}{2}}$ in the annualized version of the Sharpe. Fig.3 illustrates the very complex theoretical formula Eq.3 through a simple graph that helps to understand its sensitivity to the $\mathrm{EMA}$ parameter.
@@
In reality, Eq. 3 was derived for the one-dimensional case only but it could be easily extended to the multi-dimensional case: Obviously, Grebenkov and Serror (2015) proved that Eq. 3 holds when the correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero, as shown in their Eq. (24) [^10] Eq. (24) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}+2Q+R}}$ where $p=1-\eta$, $q=1-\lambda$, $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$, and $R=1-q^{2}-2p^{2}q^{2}$.. Eq. 3 also holds when the correlations between $\xi_{i,k}$ are the same as those between $\epsilon_{i,t}$, as shown in their Eq. (39) [^11] Eq. (39) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}+2Q+R}}$ where $p=1-\eta$, $q=1-\lambda$, $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$, and $R=1-q^{2}-2p^{2}q^{2}$., with both cases involving a scaling factor that depends only on the number of underlying assets. These latter conditions also make the Markowitz portfolio optimal, as proved by Valeyre (2024), but they are not realistic: the resulting “optimal Sharpe” would increase linearly with the square root of the number of underlyings, with the same slope regardless of correlation—an outcome that would be too ideal and profitable to be credible. This in practice, with large $N$ yields an expected Sharpe ratio based on in-sample data that is much higher—by several orders of magnitude—than the out-of-sample Sharpe ratio. This is the signature of overfitting, which occurs simply because the model assumes that the estimated means are measured without error especially for small eigenvectors, thereby allowing unrealistically profitable hedges to reduce risk under the illusion of certainty about expected mean of returns.

### v11

@@ Abstract
Our empirical results show an impressive fit with the pretty complex theoretical Sharpe formula of a trend-following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014). That empirical fit convinces us that a mean-reversion process with only one time scale is enough to model, in a pretty precise way, the reality of the trend-following mechanism at the average scale of CTAs and as a consequence, using only one simple $\mathrm{EMA}$, appears optimal to capture the trend. As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
December 31, 2025
@@ 2 Literature
The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as a signal for each underlying instrument (most often individual indicator gives either a long or a short position [^1] Sepp and Lucic (2025) classified the methodologies for sizing positions based on signals into three categories: (1) European trend-following, which assumes continuous position sizes proportional to signal strength — an approach used by many large European CTA managers; (2) Time Series Momentum (Moskowitz et al (2012); Hurst et al (2013); Baltas (2015)), a method more commonly used in academic studies, where positions depend on the sign of the momentum; and (3) American trend-following, which originates from the early adopters of CTA strategies.). Then, a risk management process is applied to the signals to ensure that the portfolio is diversified enough and targets a constant risk. One simple solution is to impose an equal conditional risk in each asset class, as the SG Trend Indicator does and to size the portfolio to target a constant volatility. Another solution is to use the correlation matrix and apply mathematical optimization to maximize the reward/risk ratio. The well-known Markowitz solution set positions as linearly depending on the signals. The linear dependence is derived simply througth the normalization of the signals by the inverse of the correlation matrix between returns, but the empirical results are not appealing. The Agnostic Risk Parity ($\mathrm{ARP}$), introduced by Benichou et al (2017), normalizes the signals through the inverse of the square root of the correlation matrix. Benichou et al (2017) proposed this portfolio because it was rotationally invariant. Since this concept is primarily understood by physicists, their portfolio is not yet widely accepted in the finance community, with only a few citations, despite posing a serious challenge to Markowitz. For example, Benveniste et al (2024) explained, for the celebration of the 50th year of the Journal of Portfolio Management, the myth surrounding the alleged inadaptation of mean–variance optimization to the real world, and how efficient mean-variance optimization works without any conflict with empirical measurements, as long as the correlation matrix is properly modeled, yet neither referenced Benichou et al (2017) nor addressed the impact of uncertainty in mean within Markowitz formula. The literature has long acknowledged the sensitivity of the Markowitz formula to parameter uncertainty, without recognizing that this uncertainty can actually alter the form of the formula itself [^2] A dedicated review of the literature is provided in the appendix..
@@
In reality, one hidden assumption of Markowitz optimization is that the expected means of returns from the signals are known and certain, but this assumption is incorrect. Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal in the context of trend-following, where expected means are not certain, are estimated with some errors and do not break rotational invariance as biases introduced by risk premia would. This optimality holds when the correlation matrix between the trends, signals or uncertainty of mean is different from the correlation matrix between returns but is a kind of random matrix where a dominant factor of this correlation matrix is very dominant and diffuses randomly. Valeyre (2024) suggested an interacting agents model to justify such an hypothesis.
@@
There are many different technical indicators used as proxies for expected means in trend-following, but all of them are based on past returns, which are supposed to help predict future ones, while inevitably containing noise and uncertainty. Most of the popular indicators used are based on moving averages of past prices. The most popular is the Simple Moving Average ($\mathrm{SMA}$), while less commonly used types include the Linear Moving Average and the Exponential Moving Average on returns ($\mathrm{EMA}$). Each moving average is computed using an averaging window of a particular size. Trend-following indicators can also be based on a combination of moving averages, such as crossovers (one with a short window size and another with a long window size). The switch is determined when the short-term moving average crosses the long-term moving average. As an example, The Societe Generale Trend Indicator, which is a reference among CTAs, is detemined only by the crossover based on a simple average based 20 and 120 business day parameters. Momentum ($\mathrm{MOM}$) is also a very popular indicator and is defined as the difference between two prices. Hurst et al (2017) analysed the performance simulated on the last century using a mixture of 1 month, 3 month and 1 year momentum. Lemperiere et al (2014) analysed the performance simulated on the last 2 centuries using a 5 month $\mathrm{EMA}$ signal [^3] They were ambiguous and only wrote “exponential moving average […] with a decay rate equal to $n$ months”. The parameter $n$ is then interpreted as the half-life, defined as the lookback period over which the cumulative weight of past observations reaches $50\%$. This corresponds approximately to $\frac{\ln(2)}{\eta}$, where $\eta$ is defined later in Eq. (6).. Gmür et al (2025) specifies that a half-life of 60-70 business days for an $\mathrm{EMA}$, corresponding to approximately one calendar quarter, best replicates the return and risk characteristics of trend-following CTA industry benchmarks such as the SG Trend Index, which invests roughly on the 10 largest systematic CTA funds with trend-following behavior, making it a robust bottom-up representation of that style. The Bollinger Band ($\mathrm{BB}$) is very popular, but non continuous with only 3 outputs 0 and 1 or -1 when price is oustide a band. The Moving Average Convergence/Divergence ($\mathrm{MACD}$) uses a combination of three $\mathrm{EMA}$s to capture mean reversion at both short and long-term scales while identifying trends in the medium term.
@@
We can describe some of these indicators through their sensitivity to past daily returns as introduced by Zakamuli and Giner (2020). Fig.1 displays some cases. It can be observed that the Exponential Moving Average ($\mathrm{EMA}$) exhibits a sensitivity that declines exponentially, while the crossover resembles a hat shape. Sensitivity is low for very recent past returns, high for a moderately distant past, and negligible for a long past. The shape of the crossover can be adapted if price behavior exhibits mean-reverting tendencies on very short time scales, as seen in the case of individual stocks with low liquidity. From an academic point of view, almost all these trend-following rules are ad hoc and lack theoretical justification for their optimality. Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33- or 36-business-day $\mathrm{EMA}$444i.e., $\eta=\frac{1}{33}$ or $\eta=\frac{1}{36}$ in Eq. (6) should be optimal for replicating the mixture of 1-month, 3-month, and 1-year momentum signals used by Hurst et al (2017), based on an equivalent average lookback or an equivalent half-life. By contrast, a 43- or 62-business-day $\mathrm{EMA}$ should be optimal for the SG Trend Indicator crossover). Zakamuli and Giner (2020) also reminds us that it is well known that if returns follow an autoregressive process, the best predictor has the same functional form as the autoregressive process being predicted. Therefore, the expected structure of the autoregressive process should provide a good proxy for the best predictor.
@@
Grebenkov and Serror (2014) proved that, in the more complex yet more insightful case where portfolio positions depend linearly on the signal [^5] Grebenkov and Serror (2014) most likely introduced into the literature the ”linear trading rule” whose position depends linearly on the indicator, rather than a purely binary trading rule, which had been the standard approach most likely because the standard approach focuses on a one-dimensional problem and on a strategy that targets a constant volatility. We were unable to find any prior published articles except Brandt and Santa Clara (2006), which is discussed in the appendix, despite consulting ChatGPT, Google Scholar, and several professors in finance. However, as we are not specialist in this specific area of the literature, it is possible that we may have missed some references. Grebenkov and Serror (2014) introduced that linear methods in the literature because I was a colleague of both and I used that method to ensure that my statistical arbitrage trading program managing more than 500 dimensions or more than 500 single stocks at John Locke Investments maintained the property of “rotational invariance.” A linear trading rule was indeed necessary to have that rotational invariance property and design consistent strategies capable of performing trend-following either on an individual stock, or a basket of stocks, or on the eigenvectors of the correlation matrix of individual stocks, rather than independently applying trend-following to each stock in isolation. , the $\mathrm{EMA}$ with a larger decay parameter than the decay parameter of the autoregressive coefficients of the return process yields the optimal portfolio. This finding is in stark disagreement with Zakamuli and Giner (2024), simply because Grebenkov and Serror (2014) solved a different, yet superior, optimization problem. Indeed, assuming that positions depend linearly on the signal appears much more natural and is not as suboptimal as merely assuming a binary choice between a long or short position with the same level of risk withtout the possibility of implementing “money management”. Most important the version of Grebenkov and Serror (2014) is in line with the ”rotational invariance” property, the Markowitz optimization and Agnostic risk parity by Benichou et al (2017) where positions are linearly depending on the signals. The linear trading rule can be extended from the one-dimensional case to the multi-dimensional setting, allowing the incorporation of the correlation matrix to better adjust portfolio positions in order to target the maximum Sharpe ratio at the portfolio level.
@@
Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant [^6] The formula proposed by Acar (2003), which depends on the correlation between the indicator and returns, leads, when expressed using the Grebenkov process, to an extremely cumbersome formula linking the parameters of the return process and the relaxation time of the $\mathrm{EMA}$, which can hardly be described as objectively elegant (description of the different formulas and the derivations when assuming the process of Grebenkov are available in the Appendix B). Furthermore, Zakamulin and Giner (2022) and Zakamuli and Giner (2024) are both based on the formula from Acar (2003). and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$. That formula is key in our main objective as we want to validate it using empirical measurements.
@@
Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community  [^7] The formula in Acar (2003); Zakamulin and Giner (2022); Zakamuli and Giner (2024) holds only with binary trading rules (proof in Appendix B.3, originaly derived by Firoozye and Koshiyama (2020)). Following Grebenkov and Serror (2014), Ferreira et al. (2018) published a similar formula for a linear trading rule, but in the context of a more general process with a simple moving average and not an $\mathrm{EMA}$. Ferreira et al. (2018) did not mention Acar (2003) and derived his results through his own way. Later, after Grebenkov and Serror (2014), Firoozye and Koshiyama (2020) extended the formula of Acar (2003) to linear trading rules, yet these developments were ignored by Zakamulin and Giner (2022) and Zakamuli and Giner (2024), who continued to focuse exclusively on binary rules while ignoring Grebenkov and Serror (2014), the literature continued to focus on one-dimensional case and the wrong optimization problem—one where positions do not depend linearly on the signals. Dai et al (2016) addressed this issue by imposing a constraint that allowed only two possible positions (long or zero). Similarly, Nguyen et al (2014b) determined the optimal threshold for triggering either a long or short position.
@@
Our first and main objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the “Agnostic Risk Portfolio” ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$. This methodology ultimately extends the idea of Ferreira et al. (2018) by focusing on an AR(1) model to capture trends and by aggregating all markets through projection onto the $\mathrm{ARP}$, rather than analyzing each market individually — which carries the risk of excessively noisy measurements and requires fitting complex and distinct ARIMA models.
@@ 3 Theory from Grebenkov and Serror (2014)
This classical one-dimensional model can be easily extended to multiple dimensions. In the multivariate case, two covariance matrices are required to account for correlations. The first matrix captures the correlations between the random components of the different underlyings, while the second matrix reflects the correlations among the AR(1) processes associated with each underlying. The model is formally described in Eq. (1), using the same notation as in Grebenkov and Serror (2014). In that equation the returns $r_{i,t}$ of the $i$th instrument among the N is described in two components: The first $\epsilon_{i,t}$ is a noise which is not autocorrelated while the second $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ is an AR(1) process which is autocorrelated where $\xi_{i,k}$ is a noise which is not autocorrelated. Ultimately Eq. (1) describes the returns of a diffusive process with a positive autocorrelation model for returns $r_{i,t}$: The weights $\left(1-\lambda\right)^{t-1-k}$ are declining exponentialy. The parameter $\lambda$ describes the inverse of the relaxation time of the mean reversion AR(1) process describing the short bias, and $\beta$ represents the strength of that short stochastic bias in the returns. The noises $\xi_{i,k}$ are independent in the $k$ axis but may be correlated with $i$ axis (Valeyre (2024) introduces specific forms of that matrix, leading to particular optimal portfolios). The noises $\epsilon_{i,t}$ are independent in the $t$ axis but may be correlated with $i$ axis with a very close correlation matrix between returns. Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ and it follows an auto-regressive model of order 1, AR(1).
@@
Grebenkov and Serror (2014) solved the correct theoretical optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals when assuming that returns are well described by Eq. (1). They considered only the one-dimensional case, but later, Grebenkov and Serror (2015) and Valeyre (2024) extended the analysis to the multi-dimensional setting.
@@
Instead of focusing on the best predictor or on the optimal portfolio with suboptimal constraints—such as the possibility of being either long with the same risk or flat, which is unfortunately common in the literature—Grebenkov and Serror (2014) assumes that positions should better depend linearly on the signals [^9] They considered only the $\mathrm{EMA}$ as the signal, which was a good choice, as Zakamuli and Giner (2024) showed that the best predictor has the same functional form as the autoregressive process being predicted. That is, in the case of Eq. (1), the optimal predictor should be an $\mathrm{EMA}$, which they indeed used., which is a natural approach which is in agreement with the Markowitz solution. They then determine the optimal signal that maximizes the Sharpe ratio. Thus, Grebenkov and Serror (2014) introduced substantial novelty, even though it is not yet widely considered by most academics specializing in trend-following.
@@
In detail, they derived an explicit formula, in Eq. (3), the theoretical Sharpe ratio, ${\mathbf{S}}\left(\eta\right)$, when assuming returns are generated through their autocorrelation model in Eq. (1), as a function of $\eta$ (the smoothing parameter of the trend indicator $\mathrm{EMA}$ which is defined mathematically later in Eq. (6) ) , $\lambda$ (the inverse of the time scale of the mean reversion process that describes the trends in Eq. (1)) and $\beta$ (the weight of the trend component in the returns Eq. (1)) or its normalized version $\beta_{0}$ while also considering the inclusion of trading costs, denoted as $\theta$. Equation Eq. (3) is in fact a remarkably elegant formula. As expected, the theoretical Sharpe ratio depends on $\beta$ (or $\beta_{0}$ its normalized version), the weight of the trend in the returns, in an approximately proportional manner when $\beta_{0}^{2}$ is significantly larger than $\frac{1}{\lambda}$ and when $\eta$ remains close to $\lambda$. In that case, the Sharpe ratio from Eq. (3) can be approximated by $\beta_{0}\sqrt{\frac{1}{2}}$ or $\beta_{0}\sqrt{\frac{255}{2}}$ in the annualized version of the Sharpe. Fig.3 illustrates the very complex theoretical formula Eq. (3) through a simple graph that helps to understand its sensitivity to the $\mathrm{EMA}$ parameter.
@@
In reality, Eq. (3) was derived for the one-dimensional case only but it could be easily extended to the multi-dimensional case: Obviously, Grebenkov and Serror (2015) proved that Eq. (3) holds when the correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero, as shown in their Eq. (24) [^10] Eq. (24) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}+2Q+R}}$ where $p=1-\eta$, $q=1-\lambda$, $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$, and $R=1-q^{2}-2p^{2}q^{2}$. Eq. (3) also holds when the correlations between $\xi_{i,k}$ are the same as those between $\epsilon_{i,t}$, as shown in their Eq. (39) [^11] Eq. (39) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}+2Q+R}}$ where $p=1-\eta$, $q=1-\lambda$, $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$, and $R=1-q^{2}-2p^{2}q^{2}$, with both cases involving a scaling factor that depends only on the number of underlying assets. These latter conditions also make the Markowitz portfolio optimal, as proved by Valeyre (2024), but they are not realistic: the resulting “optimal Sharpe” would increase linearly with the square root of the number of underlyings, with the same slope regardless of correlation—an outcome that would be too ideal and profitable to be credible. This in practice, with large $N$ yields an expected Sharpe ratio based on in-sample data that is much higher—by several orders of magnitude—than the out-of-sample Sharpe ratio. This is the signature of overfitting, which occurs simply because the model assumes that the estimated means are measured without error especially for small eigenvectors, thereby allowing unrealistically profitable hedges to reduce risk under the illusion of certainty about expected mean of returns.